• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ramesh

    Log entry time 15:34:07 on May 8,2011

    Entry number 358623

    keyword=Shift summary

    08:00 Shift starts with no beam due to RF-problem.

    Expecting 120 uA of beam for the shift and this part of the experiment is going to end at 12 AM so as to switch to SRC for the rest of the beam availability schedule.

    08:50 Injector having problem with ORO4 to deliver beam. It will take a while to fix it.

    08:52 Gas alarm for gas-shed

    09:40 Asked for controlled access to prepare for another expt: SRC. Hall will be fully surveyed. Expt folks will be in the hall by 10 AM to make hardware changes.

    10:00 Current expt ended at 09:40 so as to prepare for the next expt (SRC). After 09:40, all the day shift was used to do hardware change in the hall and do pass change from 3 to 4 pass at MCC.

    15:30 Right spectrometer polarity change from -ve to +ve finished at 14:20. Both spectrometers are set to be at correct angles (left: 20.3 deg and right: 24.5 deg). Their momentum settings are also set to be in the desired values: the left at 3.6014 GeV/c and the right at 1.19635 GeV/c. Q2,3 of both spectrometers were cycled. Right dipole is ramping up. The only work that needs to be done before we take beam is the target work which may commence in half an hour. The plan is to take beam around 18:00.

    Production runlist on 3He 20 cm at an expected 120 uA beam:

    4298 May be a junk run, mostly no beam

    4299 last run for the current expt, going to go to SRC next.